OpenStudy (anonymous):

last question: solve for each pair of eqautions x & y: x-2y=7 2x+y=34 is it the mx+b formula

OpenStudy (anonymous):

Are you supposed to be solving for x and y using this pair of equations? You could convert them both to y=mx+b, and graph them and see where they intersect, but it sounds like you're supposed to SOLVE for x & y. The first step in solving 2 equations with 2 variables is to eliminate one of the variables. You could do it with substitution: Rearrange the first equation to be x = 7 - 2y, and then plug that in for x in the second equation, so it would look like this: 2(7-2y) + y = 34 OR you could use the method where you combine the equations (can't remember what you call it off the top of my head) -- You could multiply both sides of the first equation by 2, and then subtract it from the second equation. This would eliminate the x again. 2x + y = 34 - (2x - 4y = 14) ============== -3y = 20 Either way, once you solve for y, you plug it in to the first equation, and use that info to solve for x.